ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

14. Mathematical Aspects of Computer Science

1991 MS Classification: 68, 05, 94, 90 Dusembaev, Anuar, National University of Kasakhstan, Kasakhstan An operator approach to program segmentation Relocating blocks of a program over segments of virtual memory to minimize a cost functional is known NP- complete problem. Exact solutions of this problem are obtained under very restricted conditions but most of the practical solutions are heuristic. Consider Au = f where ufromU is the segmentation task ; ffromF is the segmentation matrix and AfromH is unknown operator which is desired to be constructed for any ufromU. If H is an special algebra then the conditions for solvability of the problem of Au = f will be given. The expression of A^{∗}fromH for u^{∗}fromU^{∗}subsetU with f^{∗}fromF and the description of the subset U^{∗} will be presented. Reference: A. E. Dusembaev, One way to program segmenting, Russian Math. Doklady. 329, 6 (1993), 711-714.